Find the dimensions of a rectangle whose width is 6 miles less than its length and whose area is 55 square miles
Answer by Cromlix(4381)

Hi there,
Make the length = x
Width = x - 6
Area = length x width
55 = x(x - 6)
Rearrange
x^2 - 6x = 55
x^2 - 6x - 55 = 0
Factorize
(x - 11)(x + 5) = 0
x + 5 = 0
x = -5 disregard as -ve
x - 11 = 0
x = 11
Length = 11 miles
Width = 5 miles
Area = 11 x 5
Area = 55 miles^2
Hope this helps :-)